# Towards an Understanding of Context Utilization in Code Intelligence

### v1

@@ 7.2. Dataset
We find an uneven provision of context across different code intelligence tasks. This phenomenon is relevant to context utilization research in the corresponding task. For example, API documents are used in tasks such as code generation (Zan et al., 2023, 2022a; Ma et al., 2024a) and clone detection (Nafi et al., 2019, 2018), creating datasets that include API documents. However, many types of context remain underutilized in code intelligence tasks, relevant to conditions that datasets do not include these types of context.
Regarding programming language selection, researchers tend to create datasets for Java or Python, with other types of languages being less common. Additionally, multi-language datasets are still relatively lacking, highlighting the need for more diverse datasets to enable more accurate and comprehensive assessments. We also find that programming language selection is relevant to code intelligence tasks. For example, code generation and program repair tend to build up datasets with the context in Python language (Tang et al., 2023; Agashe et al., 2019; Yu et al., 2024; Chen et al., 2024b; Zhang et al., 2024b; Zheng et al., 2024c; Zhuo et al., 2024; Jimenez et al., 2024), while code summarization task tends to create datasets with the context in Java language (Zhou et al., 2019; Guo et al., 2023a; Wang et al., 2021a, 2023; Zhou et al., 2024; Bansal et al., 2024).
As for obtaining context, it’s relevant whether the context is direct or indirect but ultimately determined by dataset format. For example, datasets usually store source code context (Agashe et al., 2019; Yu et al., 2024; Liao et al., 2024; Zhang et al., 2024b; Zhuo et al., 2024) but less likely to store AST (Raychev et al., 2016). UML (Wang et al., 2021a) and IDE usage history (Bibaev et al., 2022) are indirect types of context but are obtainable as they are stored in datasets.
@@ 7.3. Open Source Status
Open source refers to the free availability of research data, code, model weights, and other relevant materials on public platforms. This approach promotes transparency, reproducibility, and collaboration within the research community. As shown in Table  7, we analyze the open-source status of papers in three dimensions: data, model weights, and code. “Data” refers to datasets or benchmarks created and used in the research, “model weights” denote the parameters of models proposed and trained in relevant papers, and “code” encompasses the implementation details of the models or tools described. Open source is essential for advancing research, as it enables the reproduction of results and fosters improvements through collaborative efforts.
@@
Future models are becoming increasingly complex regarding the context they process and their underlying architectures. We advocate for adopting open-source practices, encouraging researchers to make data, code, and model weights available publicly. These practices can enhance the credibility of research papers and significantly support continuous innovation.
@@ 8. Related work
The burgeoning field of code intelligence has attracted substantial research attention (Allamanis et al., 2018; Wan et al., 2024; Zheng et al., 2024a). While existing surveys predominantly concentrate on model architecture analysis, this study pioneers the first systematic review of context utilization in code intelligence, a critical yet underexplored perspective.
@@
Prior studies in other domains interpret context through domain-specific dimensions. In ubiquitous computing, Hoareau et al.  (Hoareau and Satoh, 2009) conceptualizes context through multi-dimensional attributes such as environmental states, geospatial position, and social relation. Recommender systems research (Raza and Ding, 2019; del Carmen Rodríguez-Hernández and Ilarri, 2021; Villegas et al., 2018) adopts user-centric contextual paradigms, emphasizing personal profiles and behavioral patterns. Other domains such as IoT systems (Perera et al., 2014), intelligent transportation (Huang et al., 2023), and mobile computing domains (Rivero-Rodriguez et al., 2016) also investigate the concept in their area and analyze context-aware models. Compared to these studies, our research is the first systematic review of context utilization in code intelligence tasks, as we clarify the definition of context and develop a taxonomy of contextual elements within the code intelligence domain.
@@
In the code intelligence domain, existing reviews predominantly focus on model-centric analyses. Allamanis et al. (Allamanis et al., 2018) examine machine learning models in the domain of big code, an early concept of code intelligence. They explore the similarities and differences between programming languages and natural languages, providing an in-depth review of code-relevant model designs and their applications. Yang et al.(Yang et al., 2022) review integrating deep learning techniques with software engineering, examining model selection, data processing methods, and relevant datasets. Wan et al.(Wan et al., 2024) provide a comprehensive literature review on deep learning in code intelligence, focusing on model performance and benchmarking 18 leading models across five code intelligence tasks. Zheng et al. (Zheng et al., 2024a) explore the evolution of large language models in software engineering by analyzing 134 studies, evaluating code LLMs across various code intelligence tasks, and outlining future directions for model development. In contrast, our research highlights the critical role of context in code intelligence tasks. We present a comprehensive discussion spanning rule-based, feature-based, deep learning-based, and LLM-based models across seven code intelligence tasks. We offer a more holistic perspective on how contextual information can enhance model performance and effectiveness. Moreover, we collate context-related datasets, analyze relevant metrics, and assess the current state of open-source resources in this area. This approach enables us to identify key opportunities and challenges existing studies may have overlooked, paving the way for further advancements in code intelligence.
@@ 9. Challenges and Opportunities
Our analysis identifies three core challenges in context utilization in current code intelligence systems: (1) due to the numerous potential context combinations and diverse utilization approaches, it remains unclear whether the current strategies are optimal context usage for code intelligence tasks; (2) existing methods for context processing are unexplored whether they can generalize across different code intelligence scenarios; and (3) current benchmarks and evaluation metrics cannot adequately reflect the effectiveness of context utilization.
@@
  • Opportunity 1: Integrating multiple contexts for code intelligence tasks. Different types of contexts provide distinct benefits in addressing code intelligence tasks. For example, compiler information offers detailed execution data and dynamic runtime insights, while UML diagrams present a high-level architectural view of a repository. However, existing research often focuses on a single context type or a limited combination of contexts, leaving the potential of integrating multiple contexts underexplored. While using multiple contexts may introduce additional noise, it also enriches the information available to the model. This creates both challenges and new research prospects. One promising direction is to design adaptive retrieval mechanisms that automatically adjust the types and scope of context based on task requirements and retrieval conditions, thereby improving task performance while managing computational costs.
  • Opportunity 2: Developing effective context utilization mechanisms. Handling multiple contexts can increase the complexity of models, but it also enhances their capacity to tackle more sophisticated tasks, such as repository-level code intelligence. One potential solution is to employ advanced representation learning techniques to capture semantic information more effectively. Alternatively, refining retrieval and generation strategies for relevant context can help improve context utilization. The recent advancements in large language models also offer promising opportunities for constructing more robust and scalable context-aware mechanisms.
@@
  • Opportunity 3: Constructing robust evaluations for context-aware models. Current evaluation methods for context-aware models often face challenges in adaptability when multiple contexts are introduced. Future research should focus on developing new benchmarks and evaluation metrics to address these issues. A comprehensive multidimensional framework is needed for benchmarks encompassing various code intelligence tasks and context types. For metrics, more fine-grained evaluation schemes should be explored, such as measuring the efficiency and effectiveness of contextual information processing. Additionally, enhancing model interpretability is a crucial direction. For instance, creating annotated datasets with ground-truth retrieval information would enable a more precise evaluation of context-processing capabilities.

### v2

@@ 7.2. Dataset
We find an uneven provision of context across different code intelligence tasks. This phenomenon is relevant to context utilization research in the corresponding task. For example, API documents are used in tasks such as code generation (Zan et al., 2023, 2022; Ma et al., 2024b) and clone detection (Nafi et al., 2019, 2018), creating datasets that include API documents. However, many types of context remain underutilized in code intelligence tasks, particularly for conditions that datasets do not include.
Regarding programming language selection, researchers tend to create datasets in Java or Python, with other languages less common. Additionally, multilingual datasets remain relatively scarce, underscoring the need for more diverse datasets to enable more accurate and comprehensive assessments. We also find that programming language selection is relevant to code intelligence tasks. For example, code generation and program repair tend to build up datasets with the context in Python language (Tang et al., 2023; Agashe et al., 2019; Yu et al., 2024; Chen et al., 2024; Zhang et al., 2024a; Zheng et al., 2024b; Zhuo et al., 2024; Jimenez et al., 2024), while the code summarization task tends to create datasets with the context in Java language (Zhou et al., 2019; Guo et al., 2023b; Wang et al., 2021b, 2023; Zhou et al., 2024; Bansal et al., 2024).
As for obtaining context, it’s relevant whether the context is direct or indirect but ultimately determined by dataset format. For example, datasets usually store source code context (Agashe et al., 2019; Yu et al., 2024; Liao et al., 2024; Zhang et al., 2024a; Zhuo et al., 2024) but less likely to store AST (Raychev et al., 2016). UML (Wang et al., 2021b) and IDE usage history (Bibaev et al., 2022) are indirect types of context but are obtainable as they are stored in datasets.
@@ 7.3. Open Source Status
Open source refers to the free availability of research data, code, model weights, and other relevant materials on public platforms. This approach promotes transparency, reproducibility, and collaboration within the research community. As shown in Table  8, we analyze the open-source status of papers in three dimensions: data, model weights, and code. “Data” refers to datasets or benchmarks created and used in the research, “model weights” denote the parameters of models proposed and trained in relevant papers, and “code” encompasses the implementation details of the models or tools described. Open source is essential for advancing research, as it enables the reproduction of results and fosters improvements through collaborative efforts.
@@
Future models are becoming increasingly complex regarding the context they process and their underlying architectures. We advocate adopting open-source practices, encouraging researchers to make their data, code, and model weights publicly available. These practices can enhance the credibility of research papers and significantly support continuous innovation.
@@
8. Case Study
Table 9. Performance Comparison and Context Contribution Analysis Across Code Intelligence Tasks
Task Context Paper Model Metric Benchmark Improved Method Base Performance Improved Performance Relative Improvement
Code Generation Source Code Liu et al. (Liu et al., 2023b) CodeT5 EM self-created NL+Libs+Imports(Gen)+Imports(Ret) 0.191 0.225 29.76
Code Generation API documents Zan et al. (Zan et al., 2022) CodeGen 350M Pass@1 TorchDataEval Top 1 API doc 6.72 8.72 29.76
Code Generation Compilation Information Bi et al. (Bi et al., 2024) GPT-3.5-Turbo Pass@5 CodeEval CoCoGen 34.44 43.01 24.88
Code Completion Source Code Zhang et al. (Zhang et al., 2023a) CodeGen 350M EM RepoEval RepoCoder Iteration 1 22.19 31.75 43.08
Code Completion CFG Liu et al. (Liu et al., 2024b) GPT-3.5-Turbo EM RepoEval-Updated GraphCoder 39.15 46.60 19.03
Code Completion DFG Cheng et al. (Cheng et al., 2024) CodeGen 350M EM CrossCodeEval DraCo 10.13 13.02 28.53
Code Completion Compilation Information Agrawal et al. (Agrawal et al., 2024) CodeGen 350M Compilation Rate PRAGMATICCODE CG-350M-MGD 52.43 65.37 24.69
Code Completion IDE Bibaev et al. (Bibaev et al., 2022) Decision Tree Recall@1 self-created CatBoost 0.761 0.870 14.32
Code Summarization Source Code Guo et al. (Guo et al., 2023b) LSTM-Attention METEOR self-created 14.46 MBT 16.16 11.76
Code Summarization AST Guo et al. (Guo et al., 2023b) LSTM-Attention METEOR self-created MBT&AST Analysis 16.16 16.47 1.92
Code Summarization API documents Shahbazi et al. (Shahbazi et al., 2021) Transformer METEOR CodeSearchNet API 7.60 7.82 2.89
Code Summarization Code Comments Zhou et al. (Zhou et al., 2024) CodeLlama 7B BLEU mixed project doc comments 22.50 41.00 82.22
Code Summarization UML Wang et al. (Wang et al., 2021b) GNN BLEU self-created CoCoSum 17.19 19.04 10.76
Commit Message Generation Source Code Xu et al. (Xu et al., 2022) Transformer BLEU self-created COMEG 12.12 13.47 11.14
Commit Message Generation AST Xu et al. (Xu et al., 2022) Transformer BLEU self-created COMEG 12.81 13.47 5.15
Commit Message Generation Code Diffs Vu et al. (Vu et al., 2023) UniXCoder METEOR self-created context-encoded change 5.45 6.01 10.28
Commit Message Generation Commit Messages Eliseeva et al. (Eliseeva et al., 2023) CodeT5 B-Norm CommitChronicle CodeT5: History 15.12 16.8 11.11
Commit Message Generation CPG Mandli et al. (Mandli et al., 2024) CodeT5 METEOR CodeSearchNe containing added, deleted, and context token 11.49 12.26 6.70
Defect Detection CFG Zhang et al. (Zhang et al., 2023b) CFGNN, GGNN Precision self-created NodeEmbedding+CFE 52.30 54.20 3.63
Program Repair Source Code Li et al. (Li et al., 2023b) Transformer Acr@1 DeepFix TransRepair 45.67 49.65 8.71
Program Repair Compilation Information Li et al. (Li et al., 2023b) Transformer Acr@1 DeepFix TransRepair 47.09 49.65 5.44
Program Repair Code Diffs Lin et al. (Lin et al., 2022) LSTM F1 self-created CACHE 68.30 78.00 12.40
Program Repair AST Lin et al. (Lin et al., 2022) LSTM F1 self-created CACHE 71.40 78.00 8.20
We compiled representative ablation studies that measure how different context types affect code-intelligence models. Table 9 summarizes the measured performance gains when models are provided with auxiliary context.
Each row reports a single experiment; the table’s primary focus is the relative improvement from adding context. For reference, each row also lists the task, benchmark, model, metric, context type, and the method used to obtain the gain.
The table confirms that incorporating context is a critical enhancer across all code intelligence tasks, consistently yielding non-trivial performance improvements. The observed relative improvements, ranging significantly across tasks and contexts, reveal several key insights regarding context efficacy:
  • High-Level Semantics and Documentation are Paramount. The most substantial relative improvements are often seen when leveraging external, high-level, or human-curated contexts. For example, incorporating code comments resulted in an 82.22% improvement in code summarization (Zhou et al., 2024), and using API documents led to a 29.76% gain in code generation (Zan et al., 2022). This demonstrates that domain knowledge beyond the immediate source code can yield the most dramatic increases.
  • Graph-based Context is Important for Prediction. For prediction-based tasks like Code Completion, contexts capturing program execution and data flow are crucial. The inclusion of DFG (28.53% improvement (Cheng et al., 2024)) and CFG (19.03% improvement (Liu et al., 2024b)) validates the need to model operational dependencies to achieve state-of-the-art performance explicitly.
  • Compilation Information Context Maintains Consistency. Contexts related to the development environment, such as compilation information (e.g., 24.69% for code completion (Agrawal et al., 2024)), prove vital for ensuring the practical utility and correctness of the generated or completed code.
This data strongly supports the recommendation that future code intelligence systems move beyond simple source code analysis and prioritize the integration of diverse contexts to maximize performance gains across tasks.
@@ 9. Related work
The burgeoning field of code intelligence has attracted substantial research attention (Allamanis et al., 2018; Wan et al., 2024; Zheng et al., 2024c). While existing surveys predominantly concentrate on model architecture analysis, this study pioneers the first systematic review of context utilization in code intelligence, a critical yet underexplored perspective.
@@
Prior studies in other domains interpret context through domain-specific dimensions. In ubiquitous computing, Hoareau et al.  (Hoareau and Satoh, 2009) conceptualize context through multi-dimensional attributes such as environmental states, geospatial position, and social relations. Recommender systems research (Raza and Ding, 2019; Rodríguez-Hernández2021contextSurvey; Villegas et al., 2018) adopts user-centric contextual paradigms, emphasizing personal profiles and behavioral patterns. Other domains, such as IoT systems (Perera et al., 2014), intelligent transportation (Huang et al., 2023), and mobile computing domains (Rivero-Rodriguez et al., 2016), also investigate the concept in their area and analyze context-aware models. Compared with these studies, our research is the first systematic review of context utilization in code intelligence tasks, clarifying the definition of context and developing a taxonomy of contextual elements within the code intelligence domain.
@@
In the code intelligence domain, existing reviews predominantly focus on model-centric analyses. Allamanis et al. (Allamanis et al., 2018) examine machine learning models in the domain of big code, an early concept of code intelligence. They explore the similarities and differences between programming languages and natural languages, providing an in-depth review of code-relevant model designs and their applications. Yang et al.(Yang et al., 2022) review integrating deep learning techniques with software engineering, examining model selection, data processing methods, and relevant datasets. Wan et al.(Wan et al., 2024) provide a comprehensive literature review on deep learning in code intelligence, focusing on model performance and benchmarking 18 leading models across five code intelligence tasks. Zheng et al. (Zheng et al., 2024c) explore the evolution of large language models in software engineering by analyzing 134 studies, evaluating code LLMs across various code intelligence tasks, and outlining future directions for model development. In contrast, our research highlights the critical role of context in code intelligence tasks. We present a comprehensive discussion spanning rule-based, feature-based, deep learning-based, and LLM-based models across seven code intelligence tasks. We offer a more holistic perspective on how contextual information can enhance model performance and effectiveness. Moreover, we collate context-related datasets, analyze relevant metrics, and assess the current state of open-source resources in this area. This approach enables us to identify key opportunities and challenges existing studies may have overlooked, paving the way for further advancements in code intelligence.
@@ 10. Challenges and Opportunities
Our analysis identifies three core challenges in context utilization in current code intelligence systems: (1) due to the numerous potential context combinations and diverse utilization approaches, it remains unclear whether the current strategies are optimal for context usage in code intelligence tasks. (2) existing methods for context processing are unexplored to determine whether they can generalize across different code intelligence scenarios; and (3) current benchmarks and evaluation metrics cannot adequately reflect the effectiveness of context utilization.
@@
  • Opportunity 1: Integrating multiple contexts for code intelligence tasks. Different types of contexts provide distinct benefits in addressing code intelligence tasks. For example, compiler information offers detailed execution data and dynamic runtime insights, while UML diagrams present a high-level architectural view of a repository. However, existing research often focuses on a single context type or a limited combination of contexts, leaving the potential of integrating multiple contexts underexplored. While using multiple contexts may introduce additional noise, it also enriches the information available to the model. This creates both challenges and new research prospects. One promising direction is to design adaptive retrieval mechanisms that automatically adjust the types and scope of context based on task requirements and retrieval conditions, thereby improving task performance while managing computational costs. For instance, DomainCodeBench (Zheng et al., 2025a) focuses on LLM performance across various domains with different contexts, discussing the contribution of source code context retrieved based on similarity, dependency, and API in code generation tasks. GraphCoder (Liu et al., 2024b) provides an ablation study to highlight the effects of CFG context, Data Dependency Graph (DDG) context, and Control Dependency Graph (CDG) context (the latter two are components of the PDG). While these studies have begun to offer insight, the space remains vast.
  • Opportunity 2: Developing effective context utilization mechanisms. Handling multiple contexts can increase the complexity of models, but it also enhances their capacity to tackle more sophisticated tasks, such as repository-level code intelligence. One potential solution is to employ advanced representation learning techniques to capture semantic information more effectively. Although researchers have explored representation learning methods like COMET (Mandli et al., 2024) and MGVD (Qiu et al., 2024), and achieved notable performance on related code intelligence tasks. However, the full potential of combining various context representations remains largely underexplored, and these methods often focus on performance improvements, leaving room for further investigation into time costs. Alternatively, recent advances in large language models offer promising opportunities to construct robust RAG strategies for relevant context, such as DocPrompting (Zhou et al., 2023), RLCE (Chen et al., 2024), and CoCoGen (Bi et al., 2024). A key challenge is that these frameworks often focus on specialized contexts to solve specific tasks, limiting the validation of a unified pipeline that incorporates diverse context-retrieval strategies across multiple code intelligence tasks. Moreover, future work should move beyond offline evaluation and systematically explore the timing and frequency of context extraction (e.g., offline caching and incremental updates) to bridge the gap between research prototypes and real-world production tools.
@@
  • Opportunity 3: Constructing robust evaluations for context-aware models. Current evaluation methods for context-aware models often struggle to adapt when multiple contexts are introduced. Future research should focus on developing new benchmarks and evaluation metrics to address these issues. A comprehensive, multidimensional framework is needed to benchmark various code intelligence tasks and context types. While recent research has made some initial efforts, more work is required. For example, Coderag-bench (Wang et al., 2025) provides five document sources for retrieval and is specialized for eight types of coding tasks. For metrics, more fine-grained evaluation schemes should be explored, such as measuring the efficiency and effectiveness of contextual information processing. As exemplified by RepoExec (Hai et al., 2024), which proposes the “Dependency Invocation Rate”, a metric specifically designed to quantify how well a model uses contextual information for dependency generation. Crucially, to enable the precise context evaluation, researchers need to provide more detailed ground truth annotations specific to the context when releasing benchmarks. This fine-grained assessment of context-aware metrics is an area that requires further dedicated research.
@@ 11. Conclusion and Future Work
12. Acknowledgments
This work is partially supported by CCF-Huawei Populus Grove Fund CCF-HuaweiSE202403.